# Invisible Walls in Cities: Leveraging Large Language Models to Predict Urban Segregation Experience with Social Media Content

### v2

@@ 1 Introduction
To address these challenges, we propose to leverage the reasoning power of LLMs for automated social media content analysis. We propose a Reflective LLM Coder, featuring a strategic agentic workflow that integrates two key components: a reflective attributor and a code summarizer. The reflective attributor employs an abductive reasoning approach: it prompts the LLM to estimate a location’s appeal to different demographic groups from its social media content, and reconcile discrepancies through iterative reflection on observed segregation patterns. Subsequently, the code summarizer applies chain-of-thought reasoning Wei et al. (2022) to iteratively merge insights into a structured codebook. This codebook guides LLMs to transform free-text reviews into structured summaries, highlighting segregation-related factors like cultural resonance and community engagement. Complementing this qualitative approach, we propose a REasoning and EMbedding (RE’EM) framework for quantitative segregation prediction, which combines LLMs’ reasoning capabilities with the representational power of pre-trained embedding models. First, we prompt the LLM to provide structured ratings of a place’s appeal to different groups based on the codebook, allowing the reasoning outcomes of LLMs to be vectorized and easily integrated with other channels of features. Concurrently, we fine-tune a pre-trained embedding model to learn global representations optimized for segregation prediction. Finally, the RE’EM framework fuses structured ratings, global embeddings, and population information using a neighbor-aware multi-view predictor.
We validate our approach through qualitative user studies and quantitative experiments on four US cities. In our user study with 75 researchers, participants’ prediction accuracy greatly improved when provided with LLM-generated summaries. Furthermore, as many as 80% of participants prefer the codebook-guided summaries over vanilla LLM outputs. In quantitative experiments, the RE’EM framework improves the predictive $R^{2}$ by 22.79% and reduces the MSE by 9.33% compared to baseline models relying solely on local racial composition, which is generalizable across multiple cities.
@@ 2 Related Work
Mining Web Data with LLMs. The digital footprints and rich textual contents people generate on Web platforms have proven to be informative for a wide array of social phenomena, including health outcome Nguyen et al. (2016), mental well-being Mitchell et al. (2013), crime rates Fatehkia et al. (2019), and neighborhood disparities Rama et al. (2020); Iqbal et al. (2023). However, these studies often rely on pre-calculated indices or pre-defined word lists to engineer and extract features. This is not only labor-intensive but also lacks adaptability to evolving research questions and contexts. Given the remarkable language understanding and reasoning capabilities of LLMs, recent research has explored the potential of leveraging LLMs to mine Web data for social good, such as revealing food-related social prejudice Luo et al. (2024), evaluating public accessibility Li et al. (2024), and capturing urban perception Santos et al. (2024). Nevertheless, these studies often lack an examined framework for insight extraction from massive social media content, which limits their effectiveness in capturing the full complexity of Web data. In contrast, our work presents the first step toward unlocking the reasoning power of LLMs to code free-form online reviews, extracting human-comprehensible, informative insights for segregation prediction.
@@ 3.1 Problem Formulation
Inspired by Moro et al. (2021), we compute the proportion of visitors at POI $i$ by each group $q$, denoted as $\tau_{qi}$. We then quantify the segregation at POI $i$ as the deviation of the visitor proportion $\tau_{qi}$ from the city’s residential proportion $T_{q}$ (k is a constant that normalizes $S_{i}$ to range between 0 and 1):
@@
Our task can be formulated into a POI segregation prediction problem. Given a set of POIs $I=\{i_{1},i_{2},...,i_{N}\}$ along with their corresponding social media content $C=\{c_{1},c_{2},...,c_{N}\}$ and local residential composition $L=\{l_{1},l_{2},...,l_{N}\}$, the objective is to learn a function $f(i,c,l)$ to minimize the discrepancy between the predicted segregation and the real segregation $S_{i}$:
$min~{}\text{diff}(f(i,c,l),S_{i}).$ (2)
@@ 3.2 Data
Demographic Data. We obtain the demographics of CBG residents from the American Community Survey 5-year Estimates (ACS) [^1] https://www.census.gov/programs-surveys/acs/. Specifically, we categorize the residents of CBGs into five distinct groups: hispanic, black, asian, white, and others, and calculate their respective ratios by dividing the population of each group by the total CBG population.
Social Media Data. We use the Yelp Open Dataset [^2] https://www.yelp.com/dataset for POI social media data. For each POI, we obtain its name, location, stars, categories, service attributes, and multi-modal user-generated content. The category captures its business domain, such as “food” and “department stores”. The attributes reflect its specific business traits, such as price range, delivery options, and parking availability. The user-generated content includes textual reviews and images uploaded by visitors. Each POI is associated with at least 5 reviews.
Mobility Data. To derive the ground truth of experienced segregation, we utilize the Safegraph Patterns Dataset (now accessible through Advan) [^3] https://www.deweydata.io/data-partners/advan. This dataset records monthly visits to each POI originating from various CBGs. We aggregate one year’s visitation records to construct a robust dataset with extensive observations. To address the distinct POI indexing systems between Yelp and SafeGraph, we match POIs by ensuring identical names and a location deviation within 200 meters. In Appendix A, we analyze the variability of POI features, highlighting the importance of our study.
@@ 4.1 Reflective Attributor
We design a reflective attributor to identify factors influencing a POI’s appeal to different groups. Specifically, we design a two-step Chain-of-Thought (CoT) scheme Wei et al. (2022) that guides an LLM to integrate multiple perspectives from texts and images, culminating in a reflection-driven process that refines insights based on real-world feedback.
@@
Reflection-guided Insight Refinement. Finally and arguably most importantly, we introduce real-world data to guide LLM’s reflections upon these insights. By comparing actual visitor demographics with expected residential compositions, the LLM refines the established aspects in an abductive manner, discarding inconsistencies and maintaining alignment with observed behaviors. This step not only enhances the relevance of the insights, but also provides a grounded explanation for observed visitation preferences.
@@ 4.2 Code Summarizer
Explanation Refinement. Finally, this step refines the one-sentence explanation for each code, highlighting broader, thematic insights to ensure flexibility and interpretability across contexts while preserving its analytical rigor.
After executing the whole coding process (formalized in Algorithm 1 in Appendix B), we obtain a codebook identifying 9 distinct “bricks” that form the invisible walls, shown in Table 1. Each brick represents a key aspect of the visitor experience, allowing a nuanced understanding of how places appeal to diverse racial demographics. For example, Cultural Resonance and Appeal captures how culturally themed offerings such as Italian-American cuisines attract visitors seeking authentic experiences, while Price Sensitivity and Economic Accessibility captures the impact of affordability on visitation, with moderate pricing and cost-effective policies like BYOB appealing to budget-conscious visitors.
@@ 5 REasoning-and-EMbedding (RE’EM) Framework
To effectively integrate different features for accurate segregation prediction, we design a REasoning-and-EMbedding (RE’EM) framework with four key components: a reasoning channel, a embedding channel, a population channel, and a neighbor-aware multi-view predictor, as shown in Figure 2.
@@ 5.1 Reasoning Channel
Codebook-guided Content Analysis. We instruct the LLM to analyze the given content (POI’s name, review text and images, if any) and summarize the POI’s characteristics along the 9 codebook-defined dimensions, focusing on how these may attract or repel specific racial groups.
Structured Rating. We ask the LLM to imagine itself as a member of each racial/ethnic group, and rate the POI’s attractiveness across the 9 codebook-defined dimensions based on the summary obtained in Step 1. Ratings are assigned on a scale from 0 to 10, where 0 indicates strong repulsion, 5 indicates neutrality, and 10 indicates strong attraction. This process yields 45 ratings $r_{i}$ with corresponding textual explanations for each POI $i$, providing both quantitative scores and qualitative insights into the POI’s social inclusiveness. To integrate the rating data into our predictive framework, we employ a reasoning adapter to process the LLM-generated ratings. This adapter consists of a Multi-Layer Perceptron (MLP) that transforms the ratings into a vector representation $v_{i}^{r}$, which will be subsequently fed into the predictor.
$v_{i}^{r}=\text{ReasonAdapter}(\text{LLM}(c_{i}|\text{Codebook})).$ (3)
@@ 5.2 Embedding Channel
With the embedding channel, we obtain deep representations $v_{i}^{e}$ of the review corpus for each POI $i$. Given the input token length constraint, we employ data augmentation by sampling different subsets of reviews for the same POI. We fine-tune an open-source text embedding model named GTE-base Li et al. (2023) to enhance its capability of extracting higher-level representations while preserving the pre-trained semantic knowledge. Similar to reasoning channel, the resulting embedding $e_{i}$ is adjusted to $v_{i}^{e}$ by the Embedding Adapter and then input to the predictor.
$v_{i}^{e}=\text{EmbeddingAdapter}(\text{GTE}(c_{i}))$ (4)
@@ 5.3 Population Channel
For each POI $i$, we calculate the racial composition of CBGs located within a 0.5km radius. This feature, termed $p_{i}$, provides the static demographic context based on the POI’s geographic location. We introduce the corresponding Population Adapter to obtain $v_{i}^{p}$ from $p_{i}$ for subsequent prediction.
$v_{i}^{p}=\text{PopulationAdapter}(l_{i})$ (5)
@@ 5.4 Neighbor-aware Multi-view Predictor
The neighbor-aware multi-view predictor primarily accomplishes two tasks: neighbor aggregation and multi-view fusion. For each POI, we identify the five closest neighbors in the reasoning space (similarity to $r_{i}$), the embedding space (similarity to $e_{i}$), and the geographic space (geographical proximity), respectively, forming the neighbor sets $N_{i}^{r}=\{n_{1}^{r},...,n_{5}^{r}\}$, $N_{i}^{e}=\{n_{1}^{e},...,n_{5}^{e}\}$ and $N_{i}^{p}=\{n_{1}^{p},...,n_{5}^{p}\}$. Then we employ three graph-based aggregators to aggregate the neighbors’ features $v_{n_{i}}^{r}$, $v_{n_{i}}^{e}$, and $v_{n_{i}}^{p}$:
$v_{n_{i}}^{*}=GAT_{*}(v_{i}^{*},\{v_{n}^{*}|n\in N_{i}^{*}\})$ (6)
Finally, the multi-view fusion component integrates the features from REasoning, EMbedding, and Population, combining both the individual’s $v_{i}^{*}$ and the neighbor’s $v_{n_{i}}^{*}$. This fusion process ultimately outputs the predicted segregation $\hat{S_{i}}$:
$\hat{S_{i}}=Fuse(v_{i}^{r},v_{n_{i}}^{r},v_{i}^{e},v_{n_{i}}^{e},v_{i}^{p},v_{n_{i}}^{p})$ (7)
@@ 6 User Study
We conduct a user study to evaluate the cognitive gain to humans provided by our generated codebook. The study involves 75 researchers with hands-on experience or sufficient knowledge in social media studies. To ensure that participants fully understood the survey tasks, we provided clear explanations of the geographic and racial concepts required for the study. Our questionnaire is publicly accessible  [^4] https://drive.google.com/file/d/1-NIjIPiVi_m_ONBFruYYuh4nvyhe4PF-/view?usp=sharing.
We randomly sample four POIs where visitor demographics significantly differ from local residents. For each POI, we generate review summaries using both a vanilla method (directly prompting LLMs without the codebook) and our codebook-guided method, as shown in Figure  6 in the Appendix. Participants answer four questions per POI, progressively incorporating different levels of contextual information: Baseline Estimation: Participants predict whether each racial group is over- or underrepresented at the POI based solely on its basic attributes (e.g., name, category, and service details). Review-informed Prediction: Participants reassess their estimates after being provided with a random sample of user reviews. Summary-enhanced Prediction: Participants make a final prediction after reviewing a summary of the user reviews. Summary Preference Evaluation: Participants determine whether the codebook-guided summary or the vanilla summary is more informative.
@@
Model MSE↓ Improv. RMSE↓ Improv. MAE↓ Improv. R2↑ Improv.
GTE 0.0073±0.0002 2.67% 0.0855±0.0011 1.04% 0.0685±0.0010 -0.29% 0.3249±0.0151 2.69%
BERT 0.0072±0.0002 4.00% 0.0846±0.0011 2.08% 0.0680±0.0010 0.44% 0.3394±0.0176 6.78%
GloVE 0.0072±0.0002 4.00% 0.0848±0.0011 1.85% 0.0685±0.0008 -0.29% 0.3357±0.0140 5.75%
Our Model 0.0068±0.0003 9.33% 0.0823±0.0021 4.75% 0.0662±0.0021 3.07% 0.3885±0.0278 22.79%
Table 2: Prediction performances.
@@
Figure 4 presents the results of the summary preference evaluation. Across all sampled POIs, 60%-80% of participants prefer the codebook-guided summary over the vanilla one, with an average improvement of 40%, showing that our approach more effectively distills useful information to support human judgment. As illustrated in Figure 6 in the Appendix, the codebook-guided summaries capture nuanced insights–such as potential deterrants for certain minority groups–which are often absent or less explicit without codebook guidance.

### v3

@@ 1 Introduction
To address these challenges, we propose to leverage the reasoning power of LLMs for automated social media content analysis. We propose a reflective LLM coder, featuring a strategic agentic workflow that integrates two key components: a reflective attributor and a code summarizer. The reflective attributor employs an abductive reasoning approach: it prompts the LLM to estimate a location’s appeal to different demographic groups from its social media content, and reconcile discrepancies through iterative reflection on observed segregation patterns. Subsequently, the code summarizer applies chain-of-thought reasoning Wei et al. (2022) to iteratively merge insights into a structured codebook. This codebook guides LLMs to transform free-text reviews into structured summaries, highlighting segregation-related factors like cultural resonance and community engagement. Complementing this qualitative approach, we propose a REasoning and EMbedding (RE’EM) framework for quantitative segregation prediction, which combines LLMs’ reasoning capabilities with the representational power of pre-trained embedding models. First, we prompt the LLM to provide structured ratings of a place’s appeal to different groups based on the codebook, allowing the reasoning outcomes of LLMs to be vectorized and easily integrated with other channels of features. Concurrently, we finetune an embedding model to learn global representations optimized for segregation prediction. Finally, we fuse structured ratings, global embeddings, and population information using a neighbor-aware multi-view predictor.
We validate our approach through qualitative user studies and quantitative experiments on four US cities. In our user study with 75 researchers, participants’ prediction accuracy greatly improved when provided with LLM-generated summaries. Furthermore, as many as 80% of participants prefer the codebook-guided summaries over vanilla LLM outputs. In quantitative experiments, the RE’EM framework improves the predictive $R^{2}$ by 22.79% and reduces the MSE by 9.33% compared to baseline models relying solely on local racial composition, and is generalizable across cities.
@@ 2 Related Work
Mining Web Data with LLMs. The digital footprints and rich textual contents people generate on Web platforms have proven to be informative for a wide array of social phenomena, including health outcome Nguyen et al. (2016), mental well-being Mitchell et al. (2013), crime rates Fatehkia et al. (2019), and neighborhood disparities Rama et al. (2020); Iqbal et al. (2023). However, these studies often rely on pre-calculated indices or pre-defined word lists to engineer and extract features. This is not only labor-intensive but also lacks adaptability to evolving research questions and contexts. Given the remarkable language understanding and reasoning capabilities of LLMs, recent research explored the potential of leveraging LLMs to mine Web data for social good, such as revealing food-related social prejudice Luo et al. (2024), evaluating public accessibility Li et al. (2024), and capturing urban perception Santos et al. (2024). Nevertheless, these studies often lack an examined framework for insight extraction from massive social media content, which limits their effectiveness in capturing the complexity of Web data. In contrast, our work presents the first step toward unlocking the reasoning power of LLMs to code free-form online reviews, extracting human-comprehensible, informative insights for segregation prediction.
@@ 3.1 Problem Formulation
Inspired by Moro et al. (2021), we compute the proportion of visitors at POI $i$ by each racial group $q$, denoted as $\tau_{qi}$. We then quantify the segregation at POI $i$ as the deviation of the visitor proportion $\tau_{qi}$ from the city’s residential proportion $T_{q}$ (k is a constant that normalizes $S_{i}$ to range between 0 and 1):
@@
Our task can be formulated into a POI segregation prediction problem. Given a set of POIs $I=\{i_{1},i_{2},...,i_{N}\}$ along with their corresponding social media content $C=\{c_{1},c_{2},...,c_{N}\}$ and the racial composition of local population $P=\{p_{1},p_{2},...,p_{N}\}$, the objective is to learn a function $f(i,c,p)$ to minimize the discrepancy between the predicted segregation and the real segregation $S_{i}$:
$min~\text{diff}(f(i,c,p),S_{i}).$ (2)
@@ 3.2 Data
Demographic data. We obtain the demographics of census block group (CBG) residents from the American Community Survey 5-year Estimates (ACS) [^1] https://www.census.gov/programs-surveys/acs/. A CBG is the smallest spatial unit with available demographics, typically consisting of 600-3,000 residents. We categorize CBG residents into five distinct groups: hispanic, black, asian, white, and others, and calculate their respective ratios by dividing the population of each group by the CBG population.
Social media data. We use the Yelp Open Dataset [^2] https://www.yelp.com/dataset for social media data. For each POI, we obtain its name, location, stars, categories, service attributes, and multi-modal user-generated content. The category captures its business domain, such as “food” and “department stores”. The attributes reflect its specific business traits, such as price range, delivery options, and parking availability. The user-generated content includes textual reviews and images uploaded by visitors. Each POI is associated with at least 5 reviews.
Mobility data. To derive the ground truth of experienced segregation, we utilize the Safegraph Patterns Dataset (accessed through Advan) [^3] https://www.deweydata.io/data-partners/advan. This dataset records monthly visits to each POI originating from various CBGs. We aggregate one year’s visitation records to construct a robust dataset with extensive observations. To address the distinct POI indexing systems between Yelp and SafeGraph, we match POIs by ensuring identical names and a location deviation within 200 meters. In Appendix A, we analyze the variability of POI features, highlighting the importance of our study.
@@ 4.1 Reflective Attributor
We design a reflective attributor to identify factors influencing a POI’s appeal to different groups. Specifically, we design a two-step Chain-of-Thought (CoT) scheme Wei et al. (2022) that guides an LLM to integrate multiple perspectives from texts and images, forming a reflection-driven process that refines insights based on real-world feedback.
@@
Reflection-guided insight refinement. Importantly, we introduce real-world data to guide LLM’s reflections upon these insights. By comparing actual visitor demographics with expected residential compositions, the LLM refines the established aspects in an abductive manner, discarding inconsistencies and maintaining alignment with observed behaviors. This step not only enhances the relevance of the insights, but also provides a grounded explanation for observed visitation preferences.
@@ 4.2 Code Summarizer
Explanation refinement. Finally, this step refines the one-sentence explanation for each code, highlighting broader, thematic insights to ensure flexibility and interpretability across contexts while preserving analytical rigor.
Trough the above coding process (formalized in Appendix B), we obtain a codebook identifying 9 distinct “bricks” that form the invisible walls, shown in Table 1. Each brick represents a key aspect of the visitor experience, allowing a nuanced understanding of how places appeal to diverse racial demographics. For example, Cultural Resonance and Appeal captures how culturally themed offerings such as Italian-American cuisines attract visitors seeking authentic experiences, while Price Sensitivity and Economic Accessibility captures the impact of affordability on visitation, with moderate pricing and cost-effective policies like BYOB appealing to budget-conscious visitors.
@@ 5 REasoning-and-EMbedding (RE’EM) Framework
To effectively integrate different features for accurate segregation prediction, we design a REasoning-and-EMbedding (RE’EM) framework composed of four key components: a reasoning channel, an embedding channel, a population channel, and a neighbor-aware multi-view predictor, as shown in Figure 2.
@@ 5.1 Reasoning Channel
Codebook-guided summary. For each POI $i$, we denote its associated social media content as $c_{i}$, which includes the POI’s name, review text, and images (if any). We instruct the LLM to analyze $c_{i}$ and generate a structured summary $u_{i}$ of the POI’s characteristics along the 9 codebook-defined dimensions, focusing on features that may attract or repel specific racial/ethnic groups.
Structured rating. Based on the summary $u_{i}$, we ask the LLM to imagine itself as a member of each of the 5 racial/ethnic groups and assign ratings to the POI across the 9 dimensions. Ratings range from 0 (strong repulsion) to 10 (strong attraction), with 5 indicating neutrality. This process yields a 45-dimensional vector $r_{i}$ with corresponding textual explanations $e_{i}$, providing both quantitative scores and qualitative insights into the POI’s social inclusiveness.
$r_{i},e_{i}=\text{LLM}(u_{i},\text{codebook}).$ (3)
To integrate the ratings into our predictive framework, we employ a reasoning adapter consisting of a multi-layer perceptron (MLP). It transforms the ratings into a vector representation $v_{i}^{r}$:
$v_{i}^{r}=\text{ReasonAdapter}(r_{i}).$ (4)
@@ 5.2 Embedding Channel
The embedding channel extracts deep semantic representations $v_{i}^{e}$ from the review corpus associated with each POI $i$. Due to the token length constraints, we apply data augmentation, i.e., sample different subsets of $c_{i}$ for input to the embedding model. We fine-tune an open-source text embedding model, GTE-base Li et al. (2023), to enhance its capability of extracting higher-level representations while preserving the pre-trained semantic knowledge. Although both the reasoning and embedding channels take $c_{i}$ as input, they extract complementary aspects: structured attribute reasoning in the former, and holistic semantic embedding in the latter. The resulting embedding is then passed through an embedding adapter (MLP) to produce $v_{i}^{e}$.
$v_{i}^{e}=\text{EmbeddingAdapter}(\text{GTE}(c_{i})).$ (5)
@@ 5.3 Population Channel
For each POI $i$, we compute a 5-dimensional feature vector $p_{i}$ representing the racial composition of the surrounding population. Specifically, we identify all CBGs whose centroids fall within a 0.5km radius around the POI, and compute a population-weighted average of their racial compositions. The resulting $p_{i}$ reflects the static demographic context of the POI’s geographic location. We apply a population adapter (another MLP) to transform $p_{i}$ into a latent representation $v_{i}^{p}$ for subsequent prediction.
$v_{i}^{p}=\text{PopulationAdapter}(p_{i}).$ (6)
@@ 5.4 Neighbor-Aware Multi-View Predictor
This predictor primarily performs two key operations: neighbor aggregation and multi-view fusion. For each POI $i$, we identify its five nearest neighbors in three separate spaces (views), respectively: the reasoning space (based on similarity to $r_{i}$), the embedding space (similarity to $e_{i}$), and the geographic space (physical proximity). This forms three neighbor sets $N_{i}^{r}=\{n_{1}^{r},...,n_{5}^{r}\}$, $N_{i}^{e}=\{n_{1}^{e},...,n_{5}^{e}\}$ and $N_{i}^{p}=\{n_{1}^{p},...,n_{5}^{p}\}$. We use a graph attention network (GAT) for each view to aggregate the features of the respective neighbor set, producing neighbor-aware representations $v_{n_{i}}^{r}$, $v_{n_{i}}^{e}$, and $v_{n_{i}}^{p}$:
$v_{n_{i}}^{*}=\text{GAT}_{*}(v_{i}^{*},\{v_{n}^{*}|n\in N_{i}^{*}\})$ (7)
Finally, a fully connected fusion module integrates both the POI’s representations $v_{i}^{*}$ and its neighrbors’ representations $v_{n_{i}}^{*}$ from all three channels. This fusion process ultimately outputs the predicted segregation $\hat{S_{i}}$:
$\hat{S_{i}}=\text{Fuse}(v_{i}^{r},v_{n_{i}}^{r},v_{i}^{e},v_{n_{i}}^{e},v_{i}^{p},v_{n_{i}}^{p}).$ (8)
@@ 6 User Study
We conduct a user study to evaluate the cognitive gain to humans provided by our generated codebook. The study involves 75 researchers with hands-on experience or sufficient knowledge in social media studies. To ensure that participants fully understood the survey tasks, we provided clear explanations of the geographic and racial concepts required for the study. Our questionnaire is publicly accessible  [^4] https://drive.google.com/file/d/1-NIjIPiVi_m_ONBFruYYuh4nvyhe4PF-/view.
Model MSE↓ Improv. RMSE↓ Improv. MAE↓ Improv. R2 Improv.
GTE 0.0073±0.0002 2.67%(p=0.7623) 0.0855±0.0011 1.04%(p=0.7887) 0.0685±0.0010 -0.29%(p=0.9713) 0.3249±0.0151 2.69%(p=0.8905)
BERT 0.0072±0.0002 4.00%(p=0.0147) 0.0846±0.0011 2.08%(p=0.0181) 0.0680±0.0010 0.44%(p=0.3708) 0.3394±0.0176 6.78%(p=0.0549)
GloVE 0.0072±0.0002 4.00%(p=0.0242) 0.0848±0.0011 1.85%(p=0.0290) 0.0685±0.0008 -0.29%(p=0.6118) 0.3357±0.0140 5.75%(p=0.0776)
Our Model 0.0068±0.0003 9.33%(p=0.0080) 0.0823±0.0021 4.75%(p=0.0071) 0.0662±0.0021 3.07%(p=0.0597) 0.3885±0.0278 22.79%(p=0.0022)
Table 2: Prediction performances.
We randomly sample four POIs whose visitor demographics significantly differ from local residents. For each POI, we generate review summaries using a vanilla method (directly prompting LLMs without the codebook) and our codebook-guided method (see Appendix Figure 6). Participants answer four questions per POI, progressively incorporating different levels of contextual information. (1) Baseline estimation: Participants predict whether each racial group is over- or underrepresented at the POI based solely on its basic attributes (e.g., name, category, and service details). (2) Review-informed prediction: Participants reassess their estimates after being provided with a random sample of user reviews. (3) Summary-enhanced prediction: Participants make a final prediction after reviewing a summary of the user reviews. (4) Summary preference evaluation: Participants determine whether the codebook-guided summary or the vanilla summary is more informative.
@@
Figure 4 presents the results of the summary preference evaluation. Across all sampled POIs, 60%-80% of participants prefer the codebook-guided summary over the vanilla one, with an average improvement of 40%, showing that our approach more effectively distills useful information to support human judgment. As Appendix D illustrates, the codebook-guided summaries capture nuanced insights on the potential deterrants for certain minority groups, which are often absent or less explicit without codebook guidance.